OpenStudy (anonymous):

Each of the letters in the word MATHEMATICS is represented by a letter tile in a bag. Stacey picks three letters, without replacement. To the nearest whole percent, which is the probability that she'll get all vowels?

OpenStudy (anonymous):

probability is all the possible outcomes divided by the whole set of possibilities

OpenStudy (anonymous):

how many vowels are in the word?

OpenStudy (anonymous):

\[\frac{ 4 }{ 11 }*\frac{ 3 }{ 10 }*\frac{ 2 }{ 9 }*\frac{ 1}{ 8 }\]

OpenStudy (kropot72):

\[P(all\ vowels)=\frac{4C3}{11C3}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

cuz there's no replacements

OpenStudy (kropot72):

\[\frac{4C3}{11C3}=\frac{4\times 3\times 2}{11\times 10 \times9}=you\ can\ calculate\]

OpenStudy (anonymous):

does poster have the options ?

OpenStudy (kropot72):

The question deals with sampling without replacement. In this case the hypergeometric distribution applies.

OpenStudy (anonymous):

@Luis_Rivera, Can you verify that i am right ?

OpenStudy (kropot72):

@perurabo Your calculation is for a sample of four letters, each being a vowel. The question states "picks three letters" all of the three being vowels. That is the reason I posted a correct calculation.